OpenStudy (anonymous):

what is this in simplest form... (2x-6)/(x-3)

OpenStudy (anonymous):

2(x - 6) / (x - 3) = 2(x-3)/(x-3) =2

OpenStudy (anonymous):

with what restriction?

OpenStudy (anonymous):

restriction is x cannot be 3

OpenStudy (anonymous):

since the factor gets cancelled (so I think restriction not needed)...

OpenStudy (anonymous):

oh actually there is one, since you are not allowed to cancel a zero, so \(x\) cannot be 3

OpenStudy (anonymous):

which was your first answer, stick with it